How many ways are there for four men and five women to stand in a line...

  1. How many ways are there for four men and five women to stand in a line so that:

a) All men stand together?

b) All women stand together?
Answer #1

a)here all men can be considered a unit which can be distributed within 4! ways , therefore

number of ways All men stand together = 4!*6! =24*720 =17280

b)

number of ways All women stand together =5!*5! =120*120 =14400